If current price of stock is $25 and you hold it for one year and received dividend of $2.5. You sold it at $27. How much return you received? Show dividend yield and capital gain separately.

1
Expert's answer
2021-11-28T17:48:28-0500

Dividend yield is $2.5 or "2.5\/25 = 0.1" or 10% and capital gain is 27 - 25 = $2 or "2\/25 = 0.08" or 8%.

So, the total return is 2.5 + 2 = $4.5 or 10 + 8 = 18%.